OpenStudy (asapt):

Given that MNOP is a rectangle, find x and y. A. x = 100, y= 81 B. x = 50, y = 27 C. x = 55, y = 27 D. not enough information

OpenStudy (asapt):

@ganeshie8

jimthompson5910 (jim_thompson5910):

All rectangles have 4 ninety degree angles So 2x-10 = 90 and 3y+9 = 90

OpenStudy (asapt):

so what would I do with that

jimthompson5910 (jim_thompson5910):

solve each equation for their variable

OpenStudy (asapt):

so its b?

jimthompson5910 (jim_thompson5910):

yeah you should find that 2x-10 = 90 2x = 100 x = 50

OpenStudy (asapt):

thanks can you stay on for more @jim_thompson5910

jimthompson5910 (jim_thompson5910):

sure, one more